OpenStudy (anonymous):

Solve for x to the nearest tenth. Answer A. 13.7 B. 2.1 C. 8.2 D. 5.4

OpenStudy (anonymous):



OpenStudy (anonymous):

@Calcmathlete

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

is there a website that shows some explainion for this

OpenStudy (anonymous):

Use this. \[10^2 = 9(x + 9)\]When you have a problem like this, remember that it's always \[outside \times whole = outside \times whole\]This is talking about the secant and tangent segments.

OpenStudy (anonymous):

i got 2.1

OpenStudy (anonymous):

Yup. Approximately 2.1 If you want an exact answer, it's 19/9

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

np :)